Solve each system by addition. If a unique solution does not exist, state whether the system is consistent or dependent.
2x +3y = 1
5x +3y = 16

Answer by ptaylor(1894)

Solve each system by addition. If a unique solution does not exist, state whether the system is consistent or dependent.
(1) 2x +3y = 1
(2) 5x +3y = 16
First, multiply (1) by -1 and we get:

(1) -2x -3y = -1
(2) 5x +3y = 16
Next, add (1) and (2) and we have:
3x=15
x=5
Substitute x=5 in (1):
(2)(5)+3y=1
10+3y=1 subtract 10 from both sides
3y=-10+1
3y=-9
y=-3
ck
substitute in (1)
(2)(5)+(3)(-3)=1
10-9=1
1=1
Substitute in (2)
(5)(5)+3(-3)=16
25-9=16
16=16
Hope this helps----ptaylor